# TechCoach: Towards Technical Keypoint-Aware Descriptive Action Coaching

### v1

@@ B.3 More Qualitative Results
As a supplement to Fig. 5 in the main paper, we present additional qualitative results on the generated coaching commentary in Fig. 8, further illustrating the effectiveness of our TechCoach. Moreover, Fig. 9 demonstrates that our TechCoach is capable of perceiving the action quality (i.e., Strengths and Weaknesses) of the same technical details on various action executions.
Appendix C More Details about EE4D-DAC Dataset
In this section, we show the examples of the obtained general technical keypoints, hierarchical coaching commentary, prompts for annotation collection, and other details.
@@ Appendix C More Details about EE4D-DAC Dataset
General Technical Keypoints. Completing a skilled physical action typically requires coordination across multiple body parts and objects. Thus, we define technical keypoints as language instructions spanning seven technical dimensions: 1) Head & Eyes; 2) Torso & Core; 3) Arms & Elbows; 4) Wrists & Hands; 5) Legs & Knees; 6) Feet & Heels; and 7) Human-Object Interaction. Tab. 7, Tab. 8 and Tab. 9 provide examples of the general technical keypoints for specific action tasks.
Hierarchical Coaching Commentary. Our EE4D-DAC dataset features hierarchical coaching commentary on both keypoint level and instance level. Fig. 10 and Fig. 11 show examples of the hierarchical coaching commentary.
Detailed Prompt for Annotation Collection. As described in Sec. 3.2 of the main paper, we use LLM (GPT-4o) to generate general technical keypoints and hierarchical coaching commentary. Additionally, as mentioned in Sec. 3.3, we prompt the LLM (GPT-4o) to compute the Coach-Score for commentary provided by various datasets. The detailed prompts are illustrated in Fig. 12.
@@
Manual Check on Annotations. After obtaining the annotations with LLM, we manually check the evaluation set to ensure the annotation quality.
@@ Appendix D Limitations and Future Work
Although experiments showcase the effectiveness of the proposed TechCoach, some limitations remain and warrant further exploration.
Hierarchical Technical Keypoints. In this work, for a given action task, we assume that each body part (or the dimension of human-object interaction) is associated with one paragraph of technical keypoints. While this assumption provides a general framework for the initial exploration of technical keypoint-aware DAC, technical keypoints could be further constructed hierarchically. For instance, when performing Mid-range Shooting, the shooting hand and guide hand play distinct roles in determining shooting accuracy and stability. Separating the technical keypoint for high-level body parts (e.g., wrists & hands) into finer sub-parts (e.g., shooting hand and guide hand) could further guide the model to focus on fine-grained and specific aspects in an action execution.
Pose-Assisted DAC. In this work, we focus on video-based DAC. While the Ego-Exo input enriches the visual context, there are cases where the egocentric video lacks sufficient visual information, and the human subject in the exocentric video appears relatively small. A valuable auxiliary modality to address this limitation is body pose. Since the original EgoExo4D [10] dataset provides frame-aligned 3D body pose data, incorporating pose sequences into the DAC process represents a promising direction for future research.
Figure 8: More qualitative results on the generated coaching commentary. Our TechCoach generates precise and detailed coaching commentary on what is done well and what can be improved from the given action videos. Correctly matched parts are highlighted in colors.
Figure 9: More qualitative results on the generated coaching commentary. Our TechCoach is capable of perceiving the action quality (i.e., Strengths and Weaknesses) of the same technical details (e.g., bending the knee) across various action executions. Correctly matched parts are highlighted in colors, and the corresponding regions are marked with dashed red circle. Some frames are zoomed in for better visibility.
Technical Dimensions General Technical Keypoints
@@
Table 7: General Technical Keypoints for Basketball Drills - Mikan Layup.
Technical Dimensions General Technical Keypoints
@@
Table 8: General Technical Keypoints for Rock Climbing.
Technical Dimensions General Technical Keypoints
@@
Table 9: General Technical Keypoints for Soccer Drills - Juggling.
Figure 10: An example for hierarchical coaching commentary. Zoom in for best view.
Figure 11: An example for hierarchical coaching commentary. Zoom in for best view.
Figure 12: The detailed prompts for generating General Technical Keypoints (Upper), Hierarchical Coaching Commentary (Middle), and Coach-Scores (Lower). “<Action_Domain>”: The action domain of the instance (e.g., Basketball, Soccer, Roch Climbing). “<Action_Domain>”: The action task of the instance (e.g., Reverse Layup). “<General_Technical_Keypoints>”: The generated general technical keypoints for a specific action task. “<Original_Expert_Commentary>”: The original expert commentary on the instance given by EgoExo4D [10]. Zoom in for best view.

### v2

@@ B.5 More Qualitative Results
- LLM can provide a reliable evaluation of generated text. To highlight this, we take an example from Fig. 5 in the main paper, and prompt LLM to produce accurate counts along with detailed analysis, which are shown in Fig. 9. - TechCoach is able to perceive the action quality. As a supplement to Fig. 5 in the main paper, we present additional qualitative results on the generated coaching commentary in Fig. 12, further illustrating the effectiveness of our TechCoach. Moreover, Fig. 13 demonstrates that our TechCoach is capable of perceiving the action quality (i.e., Strengths and Weaknesses) of the same technical details on various action executions.
Appendix C More Details about EE4D-DescCoach
@@
In this section, we show the examples of the obtained general TechPoints, hierarchical coaching commentary, prompts for annotation collection, and other details. - General TechPoints. Completing a skilled physical action typically requires coordination across multiple body parts and objects. Thus, we define TechPoints as language instructions spanning seven technical dimensions: (1) Head & Eyes; (2) Torso & Core; (3) Arms & Elbows; (4) Wrists & Hands; (5) Legs & Knees; (6) Feet & Heels; and (7) Human-Object Interaction. Appendix D, Appendix D and Appendix D provide examples of the general TechPoints for specific action tasks. - Hierarchical Coaching Commentary. Our EE4D-DescCoach dataset features hierarchical coaching commentary on both TechPoint level and instance level. Fig. 14 and Fig. 15 show examples of the hierarchical coaching commentary. - Detailed Prompt for Annotation Collection. As described in Sec. 3.2 of the main paper, we use LLM (GPT-4o) to generate general TechPoints and hierarchical coaching commentary. Additionally, as mentioned in Sec. 3.3, we prompt the LLM (GPT-4o) to compute the Coach-Score for commentary provided by various datasets. The detailed prompts are illustrated in Fig. 16. - Commentary length and vocabulary size. As shown in Appendix C, the vocabulary size of EE4D-DescCoach dataset is 7297. The TechPoint-level and Instance-level commentary spans average lengths of 15.8 and 68.9, respectively.
[…]
TechPoint-level 15.8$\pm$4.33 7297
Instance-level 68.9$\pm$14.1 7297
Table 9: Lengths and vocabulary size of EE4D-DescCoach dataset.
@@
Manual Check on Annotations. After obtaining the annotations with LLM, we manually check the evaluation set to ensure the annotation quality with the help of self-developed visualization tool. We filter the misaligned instances or prompt the LLM to re-generate the annotation until it meets our requirements.
@@ Appendix D Limitations and Future Work
- Scaling up TechCoach. Our TechCoach is built with a relatively small text generator (136M). We conduct fair comparisons with task-specific models training on our dataset with the same text generator and visual backbone (InternVideo2), and provide details analysis to show the contribution of the proposed Context-aware TechPoint Reasoner. However, the small text generator with well large-scale pretraining may bring shortcomings on more difficult scenarios such as out-of-domain instances. Note that action coaching in OOD scenario is still an open problem in the related field. Due to the resource and computation constraint, we currently not able to integrate an LLM (7B or larger) with our TechCoach and train the Context-aware TechPoint Reasoner from scratch. We believe scaling up TechCoach, and conducting large-scale pre- and post-training will further enhance the DescCoach performance in various scenarios. We leave them for future work. - Hierarchical TechPoints. In this work, for a given action task, we assume that each body part (or the dimension of human-object interaction) is associated with one paragraph of TechPoints. While this assumption provides a general framework for the initial exploration of TechPoint-aware DescCoach, TechPoints could be further constructed hierarchically. For instance, when performing Mid-range Shooting, the shooting hand and guide hand play distinct roles in determining shooting accuracy and stability. Separating the TechPoint for high-level body parts (e.g., wrists & hands) into finer sub-parts (e.g., shooting hand and guide hand) could further guide the model to focus on fine-grained and specific aspects in an action execution. - Pose-Assisted DescCoach. In this work, we focus on video-based DescCoach. While the Ego-Exo input enriches the visual context, there are cases where the egocentric video lacks sufficient visual information, and the human subject in the exocentric video appears relatively small. A valuable auxiliary modality to address this limitation is body pose. Since the original EgoExo4D [10] dataset provides frame-aligned 3D body pose data, incorporating pose sequences into the DescCoach process represents a promising direction for future research. Refer to caption Figure 12: More qualitative results on the generated coaching commentary. Our TechCoach generates precise and detailed coaching commentary on what is done well and what can be improved from the given action videos. Correctly matched parts are highlighted in colors. Refer to caption Figure 13: More qualitative results on the generated coaching commentary. Our TechCoach is capable of perceiving the action quality (i.e., Strengths and Weaknesses) of the same technical details (e.g., bending the knee) across various action executions. Correctly matched parts are highlighted in colors, and the corresponding regions are marked with dashed red circle. Some frames are zoomed in for better visibility.
{NiceTabular} p2.8cm|p11cm Technical Dimensions General Technical Points
@@
Table 10: General Technical Points for Basketball Drills - Mikan Layup.
{NiceTabular} p2.8cm|p11cm Technical Dimensions General Technical Points
@@
Table 11: General Technical Points for Rock Climbing.
{NiceTabular} p2.8cm|p11cm Technical Dimensions General Technical Points
@@
Table 12: General Technical Points for Soccer Drills - Juggling.
Refer to caption Figure 14: An example for hierarchical coaching commentary. Zoom in for best view.
Refer to caption Figure 15: An example for hierarchical coaching commentary. Zoom in for best view.
Refer to caption Figure 16: The detailed prompts for generating General TechPoints (Upper), Hierarchical Coaching Commentary (Middle), and Coach-Scores (Lower). “<Action_Domain>”: The action domain of the instance (e.g., Basketball, Soccer, Roch Climbing). “<Action_Domain>”: The action task of the instance (e.g., Reverse Layup). “<General_TechPoints>”: The generated general TechPoints for a specific action task. “<Original_Expert_Commentary>”: The original expert commentary on the instance given by EgoExo4D [10]. Zoom in for best view.
(a) Comparison with existing methods on score regression.